OpenStudy (anonymous):

Teach me how to write equations from drawn parabolas! The graph shows that the focus is at (0,1) and that the directrix is at y=-1.

OpenStudy (ranga):

The distance between any point on the parabola and the focus is same as the distance of that point from the directrix. You can use this fact to find the equation of the parabola. 

OpenStudy (ranga):

PF^2 = (x - 0)^2 + (y - 1)^2 PD^2 = (y - (-1))^2 PF^2 = PD^2 Plug in the values and simplify.